# Training-free Camera Control for Video Generation

### v2

@@ Abstract
We propose a training-free and robust solution to offer camera movement control for off-the-shelf video diffusion models. Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation. Instead, it can be plugged and played with most pretrained video diffusion models and generate camera controllable videos with a single image or text prompt as input. The inspiration of our work comes from the layout prior that intermediate latents hold towards generated results, thus rearranging noisy pixels in them will make output content reallocated as well. As camera move could also be seen as a kind of pixel rearrangement caused by perspective change, videos could be reorganized following specific camera motion if their noisy latents change accordingly. Established on this, we propose our method CamTrol, which enables robust camera control for video diffusion models. It is achieved by a two-stage process. First, we model image layout rearrangement through explicit camera movement in 3D point cloud space. Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images. Extensive experiments have demonstrated the robustness our method holds in controlling camera motion of generated videos. Furthermore, we show that our method can produce impressive results in generating 3D rotation videos with dynamic content. Project page at https://lifedecoder.github.io/CamTrol/.
@@ 1 Introduction
Figure 1: Training-free video camera control by CamTrol. CamTrol could handle basic camera motions, hybrid motions and complicated trajectories with precise camera coordinates. Besides, CamTrol produces impressive results on generating 3D rotation videos in various styles. Note that all human faces presented in this paper are synthesized using text-to-image models.
@@
As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension. This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes. Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions li2023videogen ; hong2022cogvideo ; ho2022imagenvideo ; luo2023videofusion ; zeng2023PixelDance ; SVD ; sora ; ge2023preserve ; fei2023DysenVDM , and demonstrated the inestimable potential of diffusion models to synthesis realistic videos. While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds zeng2023PixelDance ; SVD ; li2023videogen , most of them fail to provide camera control for the generated videos.
@@
The difficulty of controlling camera trajectory in videos primarily arises from two aspects. The initial challenge lies in the inadequacy of annotated data. Most video annotations lack of descriptions, especially precise descriptions of video’s camera movements. As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs. One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation yang2024direct . However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones. The second challenge lies in the effort of additional finetuning required for controlling camera movements. As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone. Common solutions wang2023motionctrl ; he2024cameractrl proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories. However, such datasets like RealEstate10k realestate10k and MVImageNet yu2023mvimgnet are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data. Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
@@
In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models. We begin by introducing two core observations underpin that video diffusion models can achieve camera movement control in a training-free manner. First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right. This simple implementation, though not very accurate and always leads to static or wrong motions, shows the natural prior knowledge learnt by pretrained models about following different camera trajectories. The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks. Recent works sv3d ; im3d ; shi2023mvdream find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change. The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
@@
We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner. CamTrol is inspired by the layout prior that noisy latents hold towards generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur to the output and leads to layout modification. Considering camera moves could also be seen as a kind of layout rearrangement, this prior can serve as an efficacious hint providing video model with information of specific camera motions. Specifically, CamTrol consists of a two-stage procedure. In stage I, explicit camera movements are modeled in 3D point cloud representations and produce a series of rendered images indicating specific camera movements. In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements. We conduct extensive experiments to validate the effectiveness of our proposed CamTrol. Both quantitative and qualitative results demonstrate the robustness of CamTrol as a useful tool of controlling camera motion for video diffusion models. Furthermore, we show that CamTrol  produces impressive results of dynamic 3D rotation videos in various styles.
@@ 2 Related Work
While methods aim for controlling video foundation models constantly emerge ma2024magic ; liu2023video ; feng2023ccedit , there are few works explore how to manipulate camera motion of generated videos. Earlier work hao2018controllable controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in chen2023motion ; yin2023dragnuwa . Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning. yang2024direct disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements. he2024cameractrl and wang2023motionctrl train an additional camera encoder and integrate the output into temporal attention layers of U-Net. guo2023animatediff learns new motion pattern via LoRA hu2021lora and finetuning with multiple reference videos.
@@
One of the most natural advantages of diffusion model comes from its pixel-wise noisy latents formed during denoising process. These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales. Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation meng2021sdedit , pixel-level manipulation nichol2022glide ; avrahami2022blended , image inpainting lugmayr2022repaint and semantic editing choi2021ilvr ; kim2022diffusionclip ; hou2024high . Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents mao2023guided . In other work, noise prior is used to guarantee temporal consistency among video frames luo2023videofusion , or to trade-off between fidelity and diversity of image editing kim2022diffusionclip .
@@
Similar to how most video generation models using the groundwork laid by image foundation models blattmann2023alignVideo-LDM ; esser2023GEN1 ; singer2022make ; tuneavideo , training of 3D generation model also relies heavily on pretrained 2D video models sv3d ; im3d ; shi2023mvdream ; chen2024v3d ; han2024vfusion3d . These methods either finetunes with rendered videos directly SVD ; chen2024v3d ; im3d ; han2024vfusion3d , or adds camera embedding for each view as extra condition sv3d ; shi2023mvdream . Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
@@ 3 Training-free Camera Control for Video Generation
CamTrol takes two stages to evoke the innate camera control ability hidden in base video models. In Sec. 3.1, we will describe how to model explicit camera motion in point cloud space. In Sec. 3.2, we will elaborate on the camera-controllable video generation with the guidance of noise layout prior.
@@ 3.1 Camera Motion Modeling
To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way. While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works wang2023motionctrl ; he2024cameractrl introduce additional embedder to encode camera parameters and finetune with limited annotated data realestate10k ; yu2023mvimgnet , which are data-hungry yet lack of generalization ability. Other methods yang2024direct construct camera motions by self-supervised augmentations, but could only handle a few easy camera controls. Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing explicit motion hints to video diffusion models. Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses and positions for simulating diverse camera movements. Besides explicit camera modeling, introducing point cloud brings extra benefits: The first is its data-efficiency. By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning. Second, consistency between multi-view renderings can be easily ensured, as the known points will remain unchanged once the reconstruction finished.
@@
We start by lifting the pixels in input image plane to 3D point cloud representations. In practice, the input image can be either user-defined or created by image generators like Stable Diffusion rombach2022high . Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth bhat2023zoedepth . By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
@@
where $\phi$ denotes the mapping function from RGBD to 3D point cloud, $\mathbf{K}$ and $\mathbf{P}_{0}$ represent camera’s intrinsic and extrinsic matrices set by convention chung2023luciddreamer as they’re usually intractable.
@@
To get consistent images from multiple viewpoints, we set camera motion as a pre-defined trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each of which includes a rotation matrix and translation matrix representing camera’s pose and position. At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$. By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements. Detailed definitions of these movements are elaborated in Appendix A.1. By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm. What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise coordinates, which means it can generate videos with any complicated camera motion(Fig. 1).
Multi-view Rendering
When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud. To get more reasonable results, we employ image inpainting model rombach2022high to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones. After inpainted in 2D space, image is lifted again onto 3D space and gradually complete the whole point cloud representation. During this process, misalignment between adjacent views may occur since depth estimator only estimates relative depth, further leads to inconsistency of rendered images. We adopt depth coefficient optimization chung2023luciddreamer to avoid this situation, which can be formed as:
@@ 3.2 Layout Prior of Noise
With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left[\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right]\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory. Note that quality of rendered images are not perfect as single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames. To form an ideal video, we need to find a way that satisfies the following three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
@@
Recent work on diffusion models have demonstrated the strong controllability of its noisy latents meng2021sdedit ; mao2023guided , the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models. Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image mao2023guided , so that rearranging the noise pixels will make content in output relocate as well. For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift. This reminds us that the impact of camera movement on images could also be regarded as a kind of layout rearrangement, where pixels change their positions caused by viewpoint change. In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
Inspired by this, we first construct a series of noisy latents indicating corresponding camera movements. It can be intuitively done by employing diffusion model’s inversion process on rendered image sequence $\mathbf{V}_{0}$. Latent at timestep $t_{0}$ can be calculated as:
@@
where $\bar{\alpha}_{t}$ are variances used in DDIM song2020DDIM scheduler. Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move. Moreover, while being perturbed with random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be further filled with randomness, providing video model with more possibilities to generate and correct them.
@@
Here we use DDIM as sampling scheduler song2020DDIM and $\epsilon_{\theta}$ denotes the video model for noise prediction. $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
@@
Leveraging noise prior guidance in diffusion model could lead to trade-off problem between generation’s fidelity and diversity meng2021sdedit ; kim2022diffusionclip ; hou2024high , where results that hold more faithfulness towards guidance tend to decline in generation quality or diversity. In this task, similar circumstance also occurs as model is required to receive guidance from a series of imperfect renderings while generate a reasonable and dynamic video. The key factor to balance the trade-off problem lies in the choice of $t_{0}$. When larger $t_{0}$ applied, generation bears more resemblance to original guidance $\mathbf{V}_{0}$, yet lacks of rationality and dynamics to be an appealing video. Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion. In our experiments, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance.

### v3

@@ Abstract
We propose a training-free and robust solution to offer camera movement control for off-the-shelf video diffusion models. Unlike previous work, our method does not require any supervised finetuning on camera-annotated datasets or self-supervised training via data augmentation. Instead, it can be plugged and played with most pretrained video diffusion models and generate camera controllable videos with a single image or text prompt as input. The inspiration of our work comes from the layout prior that intermediate latents hold towards generated results, thus rearranging noisy pixels in them will make output content reallocated as well. As camera move could also be seen as a kind of pixel rearrangement caused by perspective change, videos could be reorganized following specific camera motion if their noisy latents change accordingly. Established on this, we propose our method CamTrol, which enables robust camera control for video diffusion models. It is achieved by a two-stage process. First, we model image layout rearrangement through explicit camera movement in 3D point cloud space. Second, we generate videos with camera motion using layout prior of noisy latents formed by a series of rearranged images. Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment compared with other finetuned methods. Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation. Videos available at https://lifedecoder.github.io/CamTrol/.
@@ 1 Introduction
As a more appealing and content-richer modality, videos differ from images by including an extra temporal dimension. This temporal aspect provides increased versatility for depicting diverse and dynamic movements, which can be decomposed into object motion, background transitions and perspective changes. Recent years have witnessed the rapid development and splendid breakthrough of video generation with text prompt or images as input instructions (Li et al., 2023; Hong et al., 2022; Ho et al., 2022; Luo et al., 2023; Zeng et al., 2024; Blattmann et al., 2023a; Brooks et al., 2024; Ge et al., 2023; Fei et al., 2023), and demonstrated the inestimable potential of diffusion models to synthesis realistic videos. While these video generation models have made progress in generating videos with highly dynamic objects and backgrounds (Zeng et al., 2024; Blattmann et al., 2023a; Li et al., 2023), most of them fail to provide camera control for the generated videos.
@@
The difficulty of controlling camera trajectory in videos primarily arises from two aspects. The initial challenge lies in the inadequacy of annotated data. Most video annotations lack of descriptions, especially precise descriptions of video’s camera movements. As a result, video generation models trained on these data often fail to interpret text prompts related to camera motions and generate correct outputs. One solution to mitigate the data insufficiency problem is to mimic videos with camera movements through simple data augmentation (Yang et al., 2024a). However, these methods could only handle simple camera motions like zoom or truck, and have trouble in dealing with more complicated ones. The second challenge lies in the effort of additional finetuning required for controlling camera movements. As camera trajectories could be sophisticated, they sometimes cannot be accurately elaborated using naïve text prompts alone. Common solutions (Wang et al., 2023; He et al., 2024) proposed to embed camera parameters into diffusion models through learnable encoders and perform extensive finetuning on large-scale datasets with detailed camera trajectories. However, such datasets like RealEstate10k (Zhou et al., 2018) and MVImageNet (Yu et al., 2023) are intensively limited in scale and diversity due to the difficulty associated with data collection, in this way, these finetuning methods demand substantial resources but exhibit limited generalizability to other types of data. Lack of annotations and heavy finetuning effort make camera control a challenging task in video generations.
@@
In this work, we attempt to address these issues through a training-free solution to offer camera control for off-the-shelf video diffusion models. We begin by introducing two core observations underpin that video diffusion models can achieve camera movement control in a training-free manner. First, we find that base video models could produce results with rough camera moves by integrating specific camera-related text into input prompts, such as camera zooms in or camera pans right. This simple implementation, though not very accurate and always leads to static or wrong motions, shows the natural prior knowledge learnt by pretrained models about following different camera trajectories. The other observation is the effectiveness video models have exhibited in adapting to 3D generation tasks. Recent works (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023) find that leveraging pretrained video models as initialization helps drastically improve the performance of multi-view generations, demonstrating their strong ability of handling perspective change. The two crucial observations reveals the hidden power of video models for camera motion control, thus, we seek to find a way to evoke this innate ability, as it already exists in the model itself.
@@
We propose CamTrol, which offers camera control for off-the-shelf video diffusion models in a training-free but robust manner. CamTrol is inspired by the layout prior that noisy latents hold towards generation results: As pixels in noisy latents change their positions, corresponding rearrangement will also occur to the output and leads to layout modification. Considering camera moves could also be seen as a kind of layout rearrangement, this prior can serve as an efficacious hint providing video model with information of specific camera motions. Specifically, CamTrol consists of a two-stage procedure. In stage I, explicit camera movements are modeled in 3D point cloud representations and produce a series of rendered images indicating specific camera movements. In stage II, layout prior of noisy latents are utilized to guide video generations with camera movements. Compared with previous works, CamTrol require no any additional finetuning utilizing camera annotated datasets, nor does it need self-supervised training based on data augmentation. Extensive experiments have demonstrated its superior performance in both video generation quality and camera motion alignment against other finetuned methods. Furthermore, we show the capacity of CamTrol generalizing to various base models, as well as its impressive applications in scalable motion control, dealing with complicated trajectories and unsupervised 3D video generation.
@@ 2 Related Work
While methods aim for controlling video foundation models constantly emerge (Ma et al., 2024; Liu et al., 2023; Feng et al., 2023), there are few works explore how to manipulate camera motion of generated videos. Earlier work (Hao et al., 2018) controls motion trajectory via warping image through densified sparse flow and pixel fusion, similar ideas also appear later in Chen et al. (2023) and Yin et al. (2023). Besides utilizing optical flow, two main techniques for implementing video camera control are via self-supervised augmentation or additional finetuning. Yang et al. (2024a) disentangles object motion with camera movement and incorporates extra layers to embed camera motions, where model is trained in a self-supervised manner by augmenting input videos to stimulate simple camera movements. He et al. (2024) and Wang et al. (2023) train an additional camera encoder and integrate the output into temporal attention layers of U-Net. Guo et al. (2023) learns new motion pattern via LoRA (Hu et al., 2021) and finetuning with multiple reference videos.
@@
One of the most natural advantages of diffusion model comes from its pixel-wise noisy latents formed during denoising process. These latents hold strong causuality towards output and directly determine what the result looks like, meanwhile have robust error-resilience as they are perturbed by Gaussian noises across different scales. Numerous work have exploited the convenience of this noise prior to attain controllable generation, such as image-to-image translation (Meng et al., 2021), pixel-level manipulation (Nichol et al., 2022), image inpainting (Lugmayr et al., 2022) and semantic editing (Choi et al., 2021; Hou et al., 2024). Recent research has shown that even sampled from Gaussian distribution, the initial noise of diffusion process still have significantly influence to the layout of generated contents (Mao et al., 2023). In other work, noise prior is used to guarantee temporal consistency among video frames (Luo et al., 2023), or to trade-off between fidelity and diversity of image editing (Kim et al., 2022).
@@
Similar to how most video generation models using the groundwork laid by image foundation models (Blattmann et al., 2023b; Esser et al., 2023; Singer et al., 2022; Wu et al., 2023), training of 3D generation model also relies heavily on pretrained 2D video models (Voleti et al., 2024; Melas-Kyriazi et al., 2024; Shi et al., 2023; Chen et al., 2024; Han et al., 2024). These methods either finetunes with rendered videos directly (Blattmann et al., 2023a; Chen et al., 2024; Melas-Kyriazi et al., 2024; Han et al., 2024), or adds camera embedding for each view as extra condition (Voleti et al., 2024; Shi et al., 2023). Video foundation models have shown to be particularly beneficial in generating consistent multi-view rendering of 3D objects, demonstrates their inherent abundant prior knowledge for handling camera pose change.
@@ 3 Training-free Camera Control for Video Generation
CamTrol takes two stages to evoke the innate camera control ability hidden in foundation models. In Sec. 3.1, we will describe how to model explicit camera movement for video generation. In Sec. 2, we will elaborate on video’s motion control with the guidance of noise layout prior.
@@ 3.1 Camera Motion Modeling
To evoke pretrained video diffusion model’s ability of dealing with camera perspective changes, hints of camera motion should be injected to diffusion model in a proper way. While simply concatenating camera trajectories with text prompt is incomprehensible for original model, previous works (Wang et al., 2023; He et al., 2024) introduce additional embedder to encode camera parameters and finetune with limited annotated data (Zhou et al., 2018; Yu et al., 2023), which are data-hungry yet lack of generalization ability. Other method (Yang et al., 2024a) constructs camera motions by self-supervised augmentations, but could only handle a few easy camera controls. Thus, we seek a more efficient and robust way to guide the model towards camera controllable.
Considering perspective change of video is originally caused by camera movements in 3D space, we resort to 3D representation for providing generation models with explicit motion hints. Specifically, we choose point cloud as the intermediate representation, in which space we can expediently manipulate camera poses for simulating diverse movements. One extra benefit that point cloud brings is its data-efficiency: By utilizing inpainting techniques, only one single input image is required for the whole point cloud reconstruction, this sidesteps the effort of large-scale finetuning.
@@
We start by lifting the pixels in input image plane to 3D point cloud representations. In practice, the input image can be either user-defined or created by image generators like Stable Diffusion (Rombach et al., 2022). Given an input image $\mathbf{I}_{0}\in\mathbb{R}^{3\times H\times W}$, we first estimate its depth map $\mathbf{D}_{0}$ using off-the-shelf monocular depth estimator ZeoDepth (Bhat et al., 2023). By combining image and its depth map, point cloud $\mathcal{P}_{0}$ can be initialized as:
@@
where $\phi$ denotes the mapping function from RGBD to 3D point cloud, $\mathbf{K}$ and $\mathbf{P}_{0}$ represent camera’s intrinsic and extrinsic matrices set by convention (Chung et al., 2023) as they’re usually intractable.
@@
To get consistent images from multiple views, we set camera motion as a trajectory of extrinsic matrix $\left\{\mathbf{P}_{1},...,\mathbf{P}_{N-1}\right\}$, each including a rotation matrix and translation matrix representing camera’s pose and position. At each step $i$, we project the point cloud back to camera plane using function $\psi$ and get a rendered image with perspective change: $\mathbf{I}_{i}=\psi(\mathcal{P}_{i},\mathbf{K},\mathbf{P}_{i})$. By calculating extrinsic matrices of corresponding movement, we obtain a series of camera motions including zoom, tilt, pan, pedestal, truck, roll and rotate, enabling flexible camera movements. Detailed definitions of these movements are elaborated in Appendix B. By combining basic trajectories, hybrid camera movements can be attained and produce videos with cinematic charm. What’s more, benefit from explicit camera motion modeling, our method could support trajectories with precise extrinsics, which means it can generate videos with any complicated camera motion.
Multi-view Consistency
When perspective changes, there can be vacancies appear as some areas are unoccupied within the point cloud. To get more reasonable results, we employ image inpainting model (Rombach et al., 2022) to fill up the holes for new renderings, with a mask distinguishing the known points from nonexistent ones. This operation guarantees coherence between known views and novel views in 2D space. After inpainting, image is lifted again onto 3D space and gradually complete the whole point cloud representation. During this process, points between adjacent views may become misaligned since depth estimator only estimates relative depth, further leads to inconsistency in both 3D point cloud and rendered images. To avoid this situation, we adopt depth coefficient optimization (Chung et al., 2023) at each step of the camera movement, formed as:
@@ 3.2 Layout Prior of Noise
Figure 2: Transition of samples between two distinct distributions. As the layout-arranged images being inverted by adding random noise, the distribution of their noisy latents will gradually converge to their video counterpart (green area), finally forming nearly Gaussian latents with specific layout change. This information is then inherited during video generation process. Steps of camera motion inversion determine the trade-off between video diversity and motion fidelity.
With camera motion modeling, we obtain a sequence $\mathbf{V}_{0}=\left\{\mathbf{I}_{0},...,\mathbf{I}_{N-1}\right\}\in\mathbb{R}^{N\times 3\times H\times W}$ of rendered images adhering to a specific camera trajectory. Note that quality of rendered images are not perfect since single input image only leads to sparse point cloud reconstruction, besides, these renderings are static, thus they could not use directly as video frames. To form an ideal video, we need to find a way that satisfies three requirements: 1) camera motions should be maintained; 2) video should be encouraged with more dynamics; and 3) quality imperfection should be compensated.
@@
Recent work on diffusion models have demonstrated the strong controllability of its noisy latents (Meng et al., 2021; Mao et al., 2023), the causality and error-resilience they hold towards final output make them a convenient yet powerful tool for controllable generation of diffusion models. Particularly for initial noise, even sampled from Gaussian distribution, it still have significant influence on the layout of generated image (Mao et al., 2023), so that rearranging the noise pixels will make content in output relocate as well. For instance, if all pixels in initial noise shift to right by a certain distance, it is likely that generated output reflect a similar shift. This reminds us that the impact of camera movement on images could also be regarded as a kind of layout rearrangement, where pixels change their positions caused by viewpoint change. In a similar way, videos can be reorganized following camera motion if their noisy latents change accordingly.
Inspired by this, we first construct a series of noisy latents indicating specific camera movements. It can be intuitively done by employing diffusion’s inversion process on rendered image sequence $\mathbf{V}_{0}$. Latent at timestep $t_{0}$ can be calculated as follows, where $\bar{\alpha}_{t}$ are variances of the scheduler:
@@
Because the rendered images $\mathbf{V}_{0}$ share common pixels in certain regions, their latents also have relevance to each other in a way indicating pixels’ move. Moreover, while being perturbed with random noise, blank spaces and flawed regions in $\mathbf{V}_{0}$ can be further filled with randomness, providing video model with more possibilities to generate and correct them.
@@
Here $\epsilon_{\theta}$ denotes the video model for noise prediction and $\sigma_{t}$ determines whether the denoising process is deterministic or probabilistic, we set $\sigma=1$ to encourage diversity of generation results.
@@
Leveraging noise prior guidance in diffusion model could lead to a trade-off between generation’s fidelity and diversity (Meng et al., 2021; Hou et al., 2024), where results that hold more faithfulness towards guidance tend to decline in generation quality. In this task, similar circumstance also occurs as model is required to be guided by some imperfect renderings while generate a reasonable and dynamic video. The key factor to balance the trade-off problem lies in the choice of $t_{0}$. When larger $t_{0}$ applied, generation bears more resemblance to original guidance $\mathbf{V}_{0}$, yet lacks of rationality and dynamics to be an appealing video. Instead, smaller $t_{0}$ leads to well-generated video, but is less aligned with desired camera motion. Empirically, we find larger $t_{0}$ works better for motions with moderate intensity, and for those with relatively drastic move, smaller $t_{0}$ shows preferable performance. The process of stage II is illustrated in Fig. 2.